James earned 4/7 on his English quiz what was his over all percent?
She correctly answered 27 out of 45 question what was her over all percent?
I got 35 points out of 50 what was my over all percent?
Shelly's home work assignment was scored 10/25 what was her over all percent?

Answer by stanbon(75887)

James earned 4/7 on his English quiz what was his over all percent?
x/100 = 4/7
x = (4/7)100
x = 57.14%
====================
She correctly answered 27 out of 45 question what was her over all percent?
x/100 = 27/45
x = (27/45)100 = 60%
====================
I got 35 points out of 50 what was my over all percent?
x/100 = 35/50
x = 3500/50
x = 70%
------------
Shelly's home work assignment was scored 10/25 what was her over all percent?
x/100 = 10/25
x = 100(2/5)
x = 200/5
x = 40%
===============
Cheers,
Stan H.
==============